OpenStudy (abbycross167):

can someone please help me with this math question? please I'll fan and give a medal...

OpenStudy (abbycross167):

show the steps to graph this equation x - y <= -8

OpenStudy (abbycross167):

wait I'm not sure how to solve for y @Mertsj

OpenStudy (abbycross167):

so x-y<= -8-x?

OpenStudy (abbycross167):

so -1*-y>-x-8*-1?

OpenStudy (abbycross167):

@Mertsj

OpenStudy (abbycross167):

okay then what? is there another step?

OpenStudy (abbycross167):

y=x+8?

OpenStudy (abbycross167):

okay done

jimthompson5910 (jim_thompson5910):

@Mertsj the line should be solid because the inequality sign has an "or equal to" part to it

OpenStudy (abbycross167):

oh okay

OpenStudy (abbycross167):

@Mertsj all of your comments were deleted... I can't see any of them

OpenStudy (mertsj):

Jim will help you now.

OpenStudy (abbycross167):

oh okay, well thank you for helping me get a start

OpenStudy (abbycross167):

@jim_thompson5910 would you mind helping again?

jimthompson5910 (jim_thompson5910):

I'm not sure what was wrong other than the dotted line. No need to delete the comments

OpenStudy (abbycross167):

I'm not sure...

jimthompson5910 (jim_thompson5910):

where did you leave off? which part did you get to again?

OpenStudy (abbycross167):

y>x+8

jimthompson5910 (jim_thompson5910):

it's \(\Large y \ge x+8\) ? or \(\Large y > x+8\) ?

OpenStudy (abbycross167):

just > no line under it

jimthompson5910 (jim_thompson5910):

why do you have `x - y <= -8` then?

jimthompson5910 (jim_thompson5910):

<= means there's a line under the inequality

OpenStudy (abbycross167):

oh I don't know, that was just one of my answer choices... I would think my teacher would put the line under it if she meant for it to have one :/

jimthompson5910 (jim_thompson5910):

then there shouldn't be a line under the inequality in the original problem

OpenStudy (abbycross167):

so the line needs to be dashed?

jhonyy9 (jhonyy9):

<< show the steps to graph this equation x - y <= -8 >> so this mean x-y equal or less than -8 from this result y >= x+8 for x=-8 you get y=0 and for x= 1 result y= -7 - so now graph these points with coordinates (-8,0) and (1,-7)